OpenStudy (anonymous):

Solve for y in the equation 4y + 15 = –5y + 9. A. –6 B.-2/3 C.2/3 D. 6

OpenStudy (anonymous):

y = 2.667

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

lol did that too fast.

OpenStudy (anonymous):

nvm

OpenStudy (ivettef365):

combine like terms so you have 4 y + 5y = 9 -15

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i didnt see that negative

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

It's B.

OpenStudy (anonymous):

thank you Peasant1

OpenStudy (anonymous):

Sorry about the mixup. :P

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

its okay

mathslover (mathslover):

Only answers will not help you learn! you would be deprived of the joy of getting an answer on your own! It is against the \(\bf\color{red}{\href{/code-of-conduct}{Code~of~Conduct}}\) to post direct answers.

OpenStudy (anonymous):

we showed work @mathslover

OpenStudy (anonymous):

look up @mathslover

OpenStudy (anonymous):

He's just jealous we stole his spotlight. ;)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

haha

mathslover (mathslover):

The suggestion was for you all. It can be answered like this : 4y + 15 = –5y + 9. Add 5y both sides : 4y + 15 + \(\color{blue}{5y} \) = -5y + 9 + \(\color{blue}{5y}\) 9y + 15 = 9 Subtract 15 both sides 9y + 15 - \(\color{blue}{15}\) = 9 - \(\color{blue}{15}\) 9y = -6 Divide 9 both sides y = -6/9 or -2/3

OpenStudy (anonymous):

just look up we did the work easy and clearly

OpenStudy (anonymous):

and its not like they dont know how to devide or multiply or anything.

mathslover (mathslover):

It is not the matter of taking the spotlight of mine, even I am happy to see some one helping but it should be in nice manner. FriedRice, I agree that you tried to help but that's not perfect help. We ought to involve the user here too.

OpenStudy (anonymous):

ok next time or next person ill try but i wasnt trying to be rude or anything it just that i wanted to make my point.

mathslover (mathslover):

Like, you ask the user : What have you tried yet? and then the user explains what he/she has tried yet. This will make us sure that the user is not seeking of just answers. Not blaming you or anyone, but just suggesting that it would have been better if you have involved the user in conversation.

mathslover (mathslover):

Good. I am sorry, if I in any manner offended you. Thanks for considering my suggestion.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry but i dont care about medals it rediculous

OpenStudy (anonymous):

but thanks

mathslover (mathslover):

A medal shows that you're going good. Good to hear that you don't care, it is a good spirit. :)